OpenStudy (anonymous):

Can someone please explain to me how to find Total area and Lateral area of a prism? I do not understand how to use these when finding my answers.

OpenStudy (amistre64):

a prism has sides, and bases .... total area is all of it, and lateral is a word that means sides.

OpenStudy (amistre64):

lateral + bases = total area total - bases = lateral

OpenStudy (anonymous):

Thank you and how could i use that to find the T.A. of this shape?

OpenStudy (amistre64):

2 bases, and 3 sides ... what are your thoughts?

OpenStudy (anonymous):

These are formulas i have been given with my work: Lateral area= ph Total area= ph+2b Im not really certain on how i can find the perimeter, that is the part that confuses me.

OpenStudy (amistre64):

perimeter is the perimeter of a base ... perimeter is the sum of the sides of the base 

OpenStudy (amistre64):

how do we find the 3rd sides of a right triangle?

OpenStudy (amistre64):

visually that off, 6 should be bigger than 4 ... but it still works the same

OpenStudy (anonymous):

Im sorry but im not sure how..I am not very smart at this.

OpenStudy (amistre64):

one of the basic formulas you learn deals with right triangles and the relationship of their sides. something pythagorean

OpenStudy (amistre64):

its a formula that never goes away in math ... youll use it almost everywhere so you bet put it to memory a^2 + b^2 = c^2

OpenStudy (anonymous):

Oh okay, so would i do 4^2+6^2 to equal my answer..?

OpenStudy (amistre64):

close 4^2 + 6^2 = c^2 but we just want c, take the square roots sqrt(16+36) = c

OpenStudy (anonymous):

oh okay that makes more sense. So in finding that i have perimeter right? than just multiply that by height?

OpenStudy (amistre64):

yep perimeter of base is: 4+6+sqrt(52) = 10+2sqrt(13)

OpenStudy (anonymous):

Okay how would i use than with total area then?

OpenStudy (amistre64):

well you know the side area is ph and the base areas ... what is the area of a triangle?

OpenStudy (anonymous):

How would I find out the area? I'm sorry but as i have said before I am not very good with things like this.